Bowser Inc. manufacures an electronic harmonica. The following income statement shows their expected financial results at two different levels of operation.

Volume 1,000 units. 1,500 units

Sales. $100,000. $150,000

Manufacturing $50,000. $70,000
Coats

Selling and $41,000. $46,000
Administrative
Costs

Pretax Income $9,000. $34,000


What is Boweser’s Contribution Margin per unit?

What is their breakeven point in units?
Answer #1

Solution:

Variable manufacturing cost per unit = ($70,000 - $50,000) / (1500 - 1000) = $40 per unit

Fixed manufacturing cost = $50,000 - 1000*$40 = $10,000

Variable selling and administrative cost = ($46,000 - $41,000) / (1500 - 1000) = $10 per unit

Fixed selling and administrative cost = $41,000 - (1000 * $10) = $31,000

Contribution margin per unit = Selling price per unit - Variable cost per unit

= $100 - ($40 + $10) = $50 per unit

Breakeven point in units = Fixed costs / Contribution margin per unit

= ($10,000 + $31,000) / $50 = 820 units